# SolEval: Benchmarking Large Language Models for Repository-level Solidity Code Generation

### v1

@@ Abstract
Large language models (LLMs) have transformed code generation. However, most existing approaches focus on mainstream languages such as Python and Java, neglecting the Solidity language, the predominant programming language for Ethereum smart contracts. Due to the lack of adequate benchmarks for Solidity, LLMs’ ability to generate secure, cost-effective smart contracts remains unexplored. To fill this gap, we construct SolEval, the first repository-level benchmark designed for Solidity smart contract generation, to evaluate the performance of LLMs on Solidity. SolEval consists of 1,125 samples from 9 different repositories, covering 6 popular domains, providing LLMs with a comprehensive evaluation benchmark. Unlike the existing Solidity benchmark, SolEval not only includes complex function calls but also reflects the real-world complexity of the Ethereum ecosystem by incorporating gas fee and vulnerability rate. We evaluate 10 LLMs on SolEval, and our results show that the best-performing LLM achieves only 26.29% Pass@10, highlighting substantial room for improvement in Solidity code generation by LLMs. We release our data and code at https://anonymous.4open.science/r/SolEval-1C06/.
@@
Zhiyuan Peng1thanks: Equal contribution.  Xin Yin211footnotemark: 1  Rui Qian3 Peiqin Lin4,5  Yongkang Liu6  Chenhao Ying1thanks: Corresponding authors.  Yuan Luo122footnotemark: 2 1 Shanghai Jiao Tong University, China  2 Zhejiang University, China 3 Fudan University, China  4 LMU Munich, Germany 5 Munich Center for Machine Learning, Germany  6 Northeastern University, China {pzy2000, yingchenhao, yuanluo}@sjtu.edu.cnxyin@zju.edu.cn qianruii@126.comlinpq@cis.lmu.demisonsky@163.com
@@ 1 Introduction
The rapid expansion of blockchain technology and Decentralized Finance (DeFi) has led to a significant surge in smart contract deployments. This growth brings about increased development pressures and elevated security demands, highlighting the critical need for efficient and reliable Solidity code generation tools. As the cornerstone of Ethereum smart contracts, Solidity plays a fundamental role in enabling the decentralized applications that are driving the blockchain revolution.
@@
Table 1: Comparison of existing benchmarks and SolEval. Task: number of code generation tasks. SA Ratio: ratio of standalone functions. Dependency: number of dependencies (e.g., cross-file invocations). Avg. Token: average tokens in function requirements. Repo-Level: whether the benchmark is repository-level or not.
Benchmark Task SA Ratio Dependency File Avg. Token Language Repo-Level
@@
Recently, methods based on large language models (LLMs) have become the dominant approach to code generation Radford (2018); Brown et al. (2020); Yu et al. (2024). These methods can generate the corresponding functions according to descriptions in natural language. To assess the code generation capabilities of models, researchers have proposed a series of benchmarks Du et al. (2023); Yu et al. (2024); Li et al. (2024); Daspe et al. (2024). As shown in Table 1, most of these benchmarks focus on mainstream programming languages such as Python and Java, with little attention paid to the Solidity language. Different from the high flexibility of programming languages like Python, Solidity’s operation is constrained by gas fee (costs of executing operations on a blockchain) and blockchain immutability, making Solidity code generation more challenging than general programming languages. To evaluate the coding abilities of LLMs in Solidity, Daspe et al. (2024) propose the first Solidity benchmark BenchSol. However, BenchSol is entirely generated by GPT-4, distinct from real-world scenarios. Moreover, this benchmark is severely limited in scale, featuring only 15 use cases, and is restricted to evaluating LLMs on standalone functions (i.e., Non-repository-level generation).
@@
To fill the gap in Solidity benchmarks aligned with the real world, we propose SolEval, the first benchmark that supports repository-level smart contract generation. As shown in Figure 1, SolEval contains non-standalone functions that invoke context dependencies from other files, which is absent in the existing Solidity benchmark. ❶ SolEval contains 1,125 samples from 9 real-world repositories, covering 6 popular domains (e.g., security, economics, and games). ❷ SolEval is manually annotated by 5 master’s students with Solidity experience. SolEval contains detailed requirements, repositories, codes, context information, and test cases. ❸ To evaluate secure and cost-effective smart contract generation, we incorporate gas fee and vulnerability rate attributes into SolEval.
@@
We evaluate 10 popular LLMs on SolEval, including closed-source models (e.g., GPT-4o and GPT-4o-mini) and open-source models (e.g., CodeLlama and DeepSeek). The results reveal a striking performance gap: these models achieve a Pass@10 ranging from 5.91% to 26.29%, indicating that their performance in Solidity code generation is far from optimal, with significant room for improvement. The generated smart contracts exhibit varying gas fees and vulnerability rates, highlighting the dilemma of balancing cost efficiency with security in contract generation. We also have an interesting finding: DeepSeek-V3 ranks highest in Pass@10 but generates contracts with high gas fees, while DeepSeek-R1-Distill-Qwen-7B ranks lowest but generates the cheapest contracts. This contrast highlights a fundamental challenge in Solidity code generation: balancing functional correctness with gas efficiency. LLMs excelling in generating correct code may struggle with optimizing gas costs, while models focused on optimizing gas efficiency may sacrifice the quality or correctness of the generated code. Additionally, we discover that the inclusion of Retrieval-Augmented Generation (RAG) and contextual information improves model performance, highlighting the importance of incorporating contextual awareness in Solidity code generation tasks.
@@
  • We point out the limitations of the existing benchmark for Solidity smart contract generation, highlighting its insufficient scale and misalignment with real-world applications.
  • We introduce the first repository-level benchmark for Solidity smart contract generation, including a large and diverse set of 1,125 samples from 9 real-world repositories, covering 6 popular domains (i.e., security, finance, gaming, test suite, community, and gas optimization). This benchmark incorporates essential attributes such as gas fees and vulnerability rates, which are critical for smart contract development.
  • We conduct an extensive evaluation of 10 state-of-the-art LLMs on SolEval, revealing their performance gaps when generating smart contracts. We also find that LLMs can generate better contracts when using RAG and context information.
@@ 2.1 Overview
SolEval contains 1,125 samples from 9 real-world code repositories (see §A), covering 6 popular domains (e.g., security, economics, and games).
SolEval is designed for benchmarking LLMs on repository-level smart contract generation, which consists of two key phases: (1) LLM-based Solidity Code Generation (§2.2) and (2) Post-Generation Evaluation (§2.3).
2.2 LLM-based Solidity Code Generation
2.3 Post-Generation Evaluation
Following Britikov et al. (2024), we utilize Forge, which handles differences across Solidity compilers and the distribution of unit test files, to execute the test cases. We evaluate functional correctness (❻) using Pass@k and Compile@k, and assess quality attributes (❼) with Gas Fee and Vul.
Pass@k (Functional Correctness).
Pass@K measures the percentage of the problems for which at least one correctly (judged based on executing the corresponding test cases) generated solution among the top K samples generated by the LLM. To avoid the issue of high sampling variance, we use the same unbiased estimator of Pass@K implemented in HumanEval Chen et al. (2021) (see §B.3 for details).
Compile@k (Functional Compilation Correctness).
We propose the Compile@K metric to measure the percentage of the problems for which at least one correctly compiled among the top K samples generated by the LLM. Similarly to Pass@K, we count the number of samples $c^{\prime}\leq n$ that pass the compilation stage and calculate the unbiased estimator
[…]
Gas Fee (Gas Consumption).
For each sample, we use Forge to execute the corresponding test cases and calculate the gas fee, denoted as $f^{\prime}_{i}$. Then, we also calculate the gas fee of the original function from the repository, denoted as $f_{i}$. Finally, for each function sample $s$, the number of samples per function $k$, and the base LLM $l$, the intermediate gas fee is calculated by accumulating the difference $(f_{i}-f^{\prime}_{i})$ for $k$ samples per function. This result is then accumulated for all function samples $s$. Given that different LLMs can only generate the correct contract for a portion of SolEval, and that the correctly generated functions of different LLMs often do not fully intersect, we calculate gas fees only for functions in the intersection. For example, consider LLM A and LLM B: LLM A can solve problems $x$ and $y$, while LLM B can solve problems $y$ and $z$. The capabilities intersection $\mathcal{C}_{\text{intersect}}$ of LLM A and LLM B only includes problem $y$, as this is the only problem both models can handle. Thus, we restrict our gas fee calculations to the functions within this intersection, ensuring a fair comparison across the models. The total gas fee for an LLM is
Vul (Vulnerability Rate).
3 Benchmark Construction
As shown in Fig. 3, the construction of SolEval involves five key phases, each designed to ensure the robustness and diversity of the dataset. These phases are carefully structured to handle the complexities of smart contract generation, covering everything from project selection to context parsing.
Figure 3: The process of constructing SolEval.
3.1 Project Selection
To ensure SolEval’s practicality and diversity, we follow best practices Chen et al. (2021); Yu et al. (2024); Liu et al. (2024b) and select functions from different open-source projects through four steps. First, we manually select six popular GitHub organizations, such as OpenZeppelin, that host Solidity projects. We crawl all their public repositories, sort them by star count in descending order, and filter out low-star (i.e., with fewer than 40 stars) projects lacking test cases or containing fewer than 10% files written in Solidity language. By manually selecting popular GitHub projects, we ensure that SolEval assesses a model’s ability to generate smart contracts that are more likely to be used within the blockchain community.
3.2 Function Parsing
We extract all functions from the selected projects. Since native Tree-sitter Tree-sitter (2022) support for Solidity is inadequate for use, we design a Solidity version of Tree-sitter to accurately parse Solidity contracts and extract relevant information (e.g., function identifiers, bodies, and requirements). From the extracted functions, we filter out tests, interfaces, and functions with LOC smaller than five, and retain those functions invoked by test functions, successfully compiled and passed the original test cases. This process results in 1,125 function samples from different Solidity projects.
3.3 Test Construction
To further ensure the correctness of the assessment of the generated functions, we employ advanced testing techniques (i.e., Fuzz, Invariant, and Differential Testing) using Forge Foundry Book (2023). To maintain result reproducibility, we set the fuzzing seed to a fixed value (i.e., 666).
3.4 Human Annotation
We recruit five master’s students with at least three years of Solidity experience to provide double-checked, manually annotated function descriptions. There are two reasons for incorporating manually annotated comments into SolEval: (1) to reduce the LLMs’ memorization effects, as original comments are highly likely to have been encountered during the pre-training phase, and (2) to provide high-quality comments for the functions in SolEval. To ensure the quality and consistency of the annotated function descriptions, we perform an inter-annotator agreement analysis using Fleiss’ Kappa Fleiss (1971). We classify the annotated comments into four categories (i.e., intact, partially intact, unclear, and unlabeled). By calculating the observed agreement ($P_{o}$) and the expected agreement ($P_{e}$) under the assumption of independent classifications, Fleiss’ Kappa serves as a reliable indicator of annotator alignment, ranging from complete agreement ($\kappa=1$) to random agreement ($\kappa=0$). We consider $0.75\ \leq\kappa\leq 1$ an excellent level of agreement, indicating that the annotators’ decisions are highly consistent.
3.5 Context Parsing
One of the key differences between SolEval and BenchSol Daspe et al. (2024) is our consideration of contextual dependencies. In repository-level code generation, a token undefined error often occurs when the necessary context is missing, leading to compilation errors Liao et al. (2024). Therefore, providing relevant context (e.g., function signatures) is essential to help SolEval validate the model’s understanding of the requirement.
To maintain efficiency and avoid unnecessary costs or performance degradation, it is crucial to ensure that the contextual information is concise Liao et al. (2024). Following Yu et al. (2024), we define the context code (e.g., functions, variables, and interfaces) required by a function to execute as its contextual dependencies. We identify the contextual dependencies of a function through a two-step program analysis of the entire project. First, given a function to analyze, we retrieve the corresponding source file from the database, and then parse it to obtain a list of type, function, variable, and constant definitions. Next, we use static program analysis to identify all external invocations defined outside the current function, retrieving the signatures of these invocations. We then store these invocation signatures along with other relevant information about the function sample.
4 Experimental Setup
4.1 Studied LLMs
We select the 10 state-of-the-art LLMs widely used in recent code generation studies Khan et al. (2023); Yan et al. (2023); Liao et al. (2024); Yu et al. (2024); Li et al. (2024). In particular, we focus on recent models released since 2022, and we exclude the small models (with less than 1B parameters) due to their limited efficacy. Table 2 presents the 10 state-of-the-art LLMs studied in our experiments with their sizes and types. Our study includes a wide scope of LLMs that are diverse in multiple dimensions, such as (i) being both closed-source and open-source, (ii) covering a range of model sizes from 6.7B to 671B, (iii) being trained for general or code-specific purposes. For detailed descriptions of each model, refer to §B.1.
Table 2: Overview of the studied LLMs
Type Name Size
General LLM DeepSeek-V3 671B (API)
DeepSeek-R1-Distill-Qwen 7B / 32B
GPT-4o -
GPT-4o-mini -
Code LLM CodeLlama 7B / 34B
DeepSeek-Coder 6.7B / 33B
DeepSeek-Coder-V2-Lite 16B
Magicoder-S-DS 6.7B
OpenCodeInterpreter-DS 6.7B
Qwen2.5-Coder 7B / 32B
4.2 Evaluation Methodology and Metrics
We adopt the Pass@K and propose the Compile@K. The detailed explanations of the metrics are in §2.3. We set the total number (denoted as n) of samples generated by an LLM to 10, and then calculate Pass@K for the LLM with K’s value of 1, 5, and 10, respectively, which is also the case for Compile@K. When k = 1, we use the greedy search and generate a single program per requirement. When k > 1, we use the nucleus sampling with a temperature of 1 and sample k programs per requirement. We set the top-p to 0.95 and the max generation length to 512. We also use the Vul (i.e., Vulnerability Rate) and Gas Fee metrics. The detail of these metrics is illustrated in §2.3. We follow Parvez et al. (2021); Chen et al. (2024); Yin et al. (2024b) and use RAG to select the best examples and collect a database from our projects for RAG based on the functions that are excluded from SolEval. For detailed descriptions of RAG, refer to §C.3. Note that all experimental results are averaged over five independent runs.
5 Results
5.1 RQ-1 How do LLMs perform on Solidity code generation?
Table 3: Performance of LLMs on SolEval, evaluated using Pass@k, Compile@k, Gas fee (Fee), and Vulnerability Rate (Vul). The table presents results under the one-shot setting with RAG and Context. Bold values indicate the highest performance in each respective column.
LLMs Size Pass@1 Pass@5 Pass@10 Compile@1 Compile@5 Compile@10 Fee Vul
6.7B to 16B
DeepSeek-R1-Distill-Qwen 7B 2.08% 4.50% 5.91% 6.37% 18.27% 26.29% -3472 10.59%
DeepSeek-Coder-Lite 16B 10.10% 14.94% 16.79% 39.44% 54.21% 57.55% -8199 26.91%
DeepSeek-Coder 6.7B 8.39% 14.25% 16.68% 32.45% 50.74% 54.59% -7195 23.17%
CodeLlama 7B 5.15% 11.38% 14.26% 19.88% 43.05% 49.95% +18267 25.00%
Magicoder-S-DS 6.7B 7.26% 13.80% 16.68% 26.81% 48.77% 53.64% -8427 24.33%
OpenCodeInterpreter-DS 6.7B 7.05% 12.96% 15.66% 27.05% 48.71% 53.76% -8802 27.08%
Qwen2.5-Coder 7B 9.13% 15.28% 17.44% 33.31% 50.34% 54.44% -9791 29.26%
GPT-4o-mini - 7.18% 12.37% 14.69% 38.04% 53.18% 56.66% -9964 34.01%
32B to 671B
DeepSeek-V3 671B 21.72% 24.99% 26.29% 53.35% 57.57% 58.61% -7525 26.61%
DeepSeek-R1-Distill-Qwen 32B 10.19% 17.06% 19.77% 31.99% 55.31% 61.31% -7894 23.84%
DeepSeek-Coder 33B 8.32% 15.57% 18.92% 29.35% 50.08% 55.39% -8706 23.08%
CodeLlama 34B 6.80% 13.52% 16.47% 24.59% 48.68% 54.80% -8412 25.47%
Qwen2.5-Coder 32B 13.46% 19.28% 21.44% 44.03% 55.53% 57.87% -7959 24.52%
GPT-4o - 12.96% 20.79% 23.70% 47.04% 58.45% 60.74% -9640 21.50%
Evaluation of Pass@k and Compile@k for generated code. Table 3 presents the overall performance of state-of-the-art LLMs on SolEval. Among the 6.7B-to-16B models, DeepSeek-Coder-Lite achieves the highest Pass@1 and Compile@1, surpassing other models. Notably, DeepSeek-R1-Distill-Qwen-7B, which claims comparable performance to ChatGPT-o1-mini on benchmarks such as LiveCodeBench and CodeForces DeepSeek (2025), underperforms compared to CodeLlama-7B. This discrepancy is likely due to DeepSeek-R1-Distill’s lack of knowledge of Solidity, highlighting the importance of a specialized benchmark like SolEval. Among the 32B-to-34B models, Qwen2.5-Coder outperforms others in both Pass@k and Compile@k. Overall, DeepSeek-V3 performs best with a 26.29% Pass@10. It is noteworthy that DeepSeek-R1-Distill-Qwen-32B significantly outperforms its 7B counterpart, maintaining most of its Solidity code generation capabilities.
Evaluation of Gas Fee and Vulnerability Rate for generated code. As shown in Table 3, there is a significant variation in gas fee and vulnerability rate across various LLMs. DeepSeek-V3 ranks first in Pass@k but generates the most gas-inefficient contracts among the 32B-to-671B models. Additionally, GPT-4o-mini, while being outperformed by GPT-4o in Pass@k and vulnerability rate, excels in generating contracts with lower gas fee.
Impact of different example numbers. As previous studies Brown et al. (2020); Liao et al. (2024) have shown, the number of examples provided has a significant impact on LLMs’ performance. To explore this, we adjust the number of examples while keeping other parameters and hyperparameters constant to ensure a fair comparison. We do not conduct experiments in a zero-shot setting, as LLMs may generate unnormalized outputs without a prompt template, which would hinder automated extraction. From Fig. 4, we observe that as the number of examples increases, both the average token length and time cost rise sharply, while the improvement in Pass@k remains modest. Based on these findings, we perform our ablation studies (Table 3 and 4) using a one-shot setting in SolEval.
Impact of Context Information. Since that relevant context typically enhances performance in other programming languages, we conduct an ablation study to examine the influence of context on the quality of LLM-generated contracts. Table 4 shows that providing context information improves both Pass@1 and Compile@1. However, there is no clear correlation between gas fees, vulnerability rate, and the presence of context information.
RAG Context Pass@1 Compile@1 Fee Vul
21.72% 53.35% -7525 26.61%
20.24% 51.08% 3828 23.68%
21.28% 52.54% -708 26.13%
20.17% 50.32% 768 26.83%
5.3 Empirical Lessons
Larger language models do not necessarily improve the gas fee of the generated code.
In Table 3, DeepSeek-V3 (671B) ranks first in Pass@k but generates the most gas-inefficient contracts among the 32B-to-671B LLMs. Furthermore, GPT-4o-mini is outperformed by GPT-4o in Pass@k but excels in crafting contracts that cost less gas fee.
6 Related Work
6.1 Large Language Model
The advancement of pre-training technology has significantly advanced code generation in both academia and industry  Li et al. (2022); Shen et al. (2022); Nijkamp et al. (2022); Fried et al. (2023). This has led to the emergence of numerous Large Language Models (LLMs) that have made substantial strides in code generation, including ChatGPT OpenAI (2022), Magicoder Wei et al. (2023), CodeLlama Roziere et al. (2023), and Qwen Bai et al. (2023), DeepSeek-Coder DeepSeek (2024b) and OpenCodeInterpreter Zheng et al. (2024).
To optimize LLMs for various code generation scenarios, some previous studies focus on enhancing prompt engineering by introducing specific patterns, such as Structured Chain-of-Thought Yin et al. (2024b); Li et al. (2025), Self-planning Jiang et al. (2024b), Self-debug Chen et al. (2023); Xia and Zhang (2023), and Self-collaboration Dong et al. (2024); Yin et al. (2024a). However, these efforts primarily address mainstream programming languages (e.g., Java, Python, and C++) Yin et al. (2024a, c); Xia and Zhang (2023).
6.2 Code Generation Benchmark
7 Conclusion and Future Work
This paper presents a new benchmark named SolEval to evaluate LLMs’ effectiveness in Solidity smart contract generation scenarios. Compared with BenchSol Daspe et al. (2024), SolEval supports repository-level smart contract generation and excels in scale (75 times in number of tasks) and real-world code alignment. Meanwhile, our benchmark takes vulnerability rate and gas fee into consideration, both of which are crucial for secure and cost-effective smart contract development. The experimental results show that SolEval can reveal the weaknesses of 10 state-of-the-art LLMs, highlighting the limitations of these LLMs in generating non-standalone Solidity functions.
In the future, there are two main directions for extending SolEval. Firstly, we will look for more high-quality code repositories from GitHub and enlarge our dataset with more projects and test cases. Secondly, we plan to support more programming languages to make it a multilingual benchmark.
Limitations
We believe that SolEval has four limitations:
  • Due to funding constraints, we were unable to evaluate SolEval on GPT-o3-mini-high and its competitors (e.g., Claude 3.5) in our study. This limitation may affect the generalizability of our findings, as these models have demonstrated advanced capabilities in various benchmarks.
  • The function samples in SolEval are drawn from 9 GitHub repositories, which may not be sufficient for a benchmark on par with those for mainstream programming languages. However, given the limited availability and accessibility of high-quality Solidity datasets, we have made a trade-off between repository quality and quantity.
  • The gas fee and vulnerability rate metrics used in SolEval are limited to evaluating the gas efficiency and potential vulnerabilities of smart contracts without providing mechanisms for their optimization or remediation. In future work, we plan to extend our research to include methods for gas optimization and vulnerability detection, thereby enhancing the practical applicability of SolEval in improving smart contract performance and security.
Ethics Consideration
Appendix A Statistics of SolEval
The statistics for the 9 projects are shown in Table 5. The functions that are filtered out can still serve as knowledge databases for RAG to select examples.
Table 5: The statistics of the 9 projects. Fi.: Filtered Functions with filter rules defined in Section 3.1.
Project Function Test Case LOC

### v3

@@ Abstract
Large language models (LLMs) have transformed code generation. However, most existing approaches focus on mainstream languages such as Python and Java, neglecting the Solidity language, the predominant programming language for Ethereum smart contracts. Due to the lack of adequate benchmarks for Solidity, LLMs’ ability to generate secure, cost-effective smart contracts remains unexplored. To fill this gap, we construct SolEval, the first repository-level benchmark designed for Solidity smart contract generation, to evaluate the performance of LLMs on Solidity. SolEval consists of 1,507 samples from 28 different repositories, covering 6 popular domains, providing LLMs with a comprehensive evaluation benchmark. Unlike the existing Solidity benchmark, SolEval not only includes complex function calls but also reflects the real-world complexity of the Ethereum ecosystem by incorporating Gas@k and Vul@k. We evaluate 16 LLMs on SolEval, and our results show that the best-performing LLM achieves only 26.29% Pass@10, highlighting substantial room for improvement in Solidity code generation by LLMs. Additionally, we conduct supervised fine-tuning (SFT) on Qwen-7B using SolEval, resulting in a significant performance improvement, with Pass@5 increasing from 16.67% to 58.33%, demonstrating the effectiveness of fine-tuning LLMs on our benchmark. We release our data and code at https://github.com/pzy2000/SolEval.
@@
Zhiyuan Peng1,2thanks: Equal contribution.  Xin Yin311footnotemark: 1  Rui Qian4  Peiqin Lin5,6 Yongkang Liu7  Hao Zhang8  Chenhao Ying1thanks: Corresponding authors.  Yuan Luo122footnotemark: 2 1 Shanghai Jiao Tong University, China  2 Shanghai SimMed Technology Co, Ltd 3 Zhejiang University, China  4 Fudan University, China 5 LMU Munich, Germany  6 Munich Center for Machine Learning, Germany 7 Northeastern University, China  8 Universiti Sains Malaysia, Malaysia {pzy2000,yingchenhao,yuanluo}@sjtu.edu.cn xyin@zju.edu.cnqianruii@126.com linpq@cis.lmu.demisonsky@163.comzhanghao666@student.usm.my
@@ 1 Introduction
The rapid expansion of blockchain technology and Decentralized Finance (DeFi) has led to a significant surge in smart contract deployments. This growth brings about increased development pressures and elevated security demands, highlighting the critical need for efficient and reliable Solidity code generation tools Qian et al. (2023); Chaliasos et al. (2024). As the cornerstone of Ethereum contracts, Solidity plays a fundamental role in enabling the decentralized applications that are driving the blockchain revolution Smaragdakis et al. (2025).
@@
Table 1: Comparison of existing benchmarks and SolEval. Sample: number of class/function samples. SA Ratio: ratio of standalone functions. Dependency: number of dependencies (e.g., cross-file invocations). Avg. Token: average tokens in function requirements. Repo-Level: whether the benchmark is repository-level or not.
Benchmark Sample SA Ratio Dependency File Avg. Token Language Repo-Level
@@
Recently, methods based on large language models (LLMs) have become the dominant approach to code generation Radford (2018); Brown et al. (2020); Yu et al. (2024). These methods can generate the corresponding functions according to descriptions in natural language. To assess the code generation capabilities of models, researchers have proposed a series of benchmarks Du et al. (2023); Yu et al. (2024); Li et al. (2024); Daspe et al. (2024). As shown in Table 1, most of these benchmarks focus on mainstream programming languages such as Python and Java, with little attention paid to Solidity. Different from the high flexibility of programming languages like Python, Solidity’s operation is constrained by gas fee (costs of executing operations on a blockchain) and blockchain immutability, making Solidity code generation more challenging than general programming languages. To evaluate the coding abilities of LLMs in Solidity, Daspe et al. (2024) proposes the first Solidity benchmark, BenchSol. However, BenchSol is entirely generated by GPT-4, distinct from real-world scenarios. Moreover, this benchmark is severely limited in scale, featuring only 15 functions, and is restricted to standalone functions (i.e., Non-repository-level generation). Cross-file calls, library imports, and storage layout, which are absent in Non-repository-level benchmarks, directly reflect the complexity of real-world smart contract development.
@@
To fill the gap, we propose SolEval, the first benchmark that supports repository-level smart contract generation. As shown in Figure 1, SolEval contains non-standalone functions that invoke context dependencies from other files, which are absent in the existing Solidity benchmark. ❶ SolEval contains 1,507 samples from 28 real-world repositories, covering 6 popular domains (e.g., security, economics, and games). ❷ SolEval is manually annotated by 5 master’s students with Solidity experience. SolEval contains detailed requirements, repositories, codes, context information, and test cases. ❸ To evaluate secure and cost-effective smart contract generation, we incorporate Gas@k and Vul@k attributes into SolEval.
@@
We evaluate 16 popular LLMs on SolEval, including closed-source models (e.g., GPT-4o and GPT-4o-mini) and open-source models (e.g., CodeLlama and DeepSeek-R1). The results reveal a striking performance gap: these models achieve a Pass@10 ranging from 5.91% to 26.29%, indicating that their performance in Solidity code generation is far from optimal, with significant room for improvement. The generated smart contracts exhibit varying gas fees and vulnerability rates, highlighting the dilemma of balancing cost efficiency with security in contract generation.
We also have an interesting finding: DeepSeek-V3 ranks highest in Pass@10 but generates contracts with high gas fees, while DeepSeek-R1-Distill-Qwen-7B ranks lowest but generates the cheapest contracts. This contrast highlights a fundamental challenge in Solidity code generation: balancing functional correctness with gas efficiency. LLMs excelling in generating correct code may struggle with optimizing gas costs, while models focused on optimizing gas efficiency may sacrifice the quality or correctness of the generated code.
Additionally, we discover that the inclusion of Retrieval-Augmented Generation (RAG) and contextual information improves model performance, highlighting the importance of incorporating contextual awareness in Solidity code generation tasks. In particular, we conduct supervised fine-tuning (SFT) on Qwen-7B using SolEval, resulting in a significant performance improvement. Pass@5 increases from 16.67% to 58.33%, demonstrating that fine-tuning LLMs on our benchmark leads to a notable enhancement in the generation of high-quality Solidity code. This reinforces the effectiveness of our benchmark in improving LLM performance through task-specific training.
@@
  • We introduce the first repository-level benchmark for Solidity code generation, including a diverse set of 1,507 samples from 28 real-world repositories, covering 6 popular domains. We also propose essential metrics (i.e., Gas@k and Vul@k) critical for smart contract development.
  • We conduct an extensive evaluation of 16 state-of-the-art LLMs on SolEval, revealing their performance gaps when generating smart contracts. We find that LLMs can generate better contracts when using RAG and context information.
  • We conduct supervised fine-tuning (SFT) on Qwen-7B using SolEval, demonstrating a significant performance improvement, with Pass@5 increasing from 16.67% to 58.33%. This substantial enhancement highlights the effectiveness of fine-tuning LLMs on our benchmark for generating high-quality Solidity smart contracts.
@@ 2.1 Overview
SolEval contains 1,507 samples from 28 real-world code repositories, covering 6 popular domains (e.g., security, economics, and games). The statistics for the projects are shown in Table 2. The functions that are filtered out can still serve as knowledge databases for RAG to select examples.
Table 2: The simplified statistics of the top-9 projects. Fi.: Filtered Functions with rules defined in Section 3.1.
Top 9 Project Function Test Case LOC